Suppose we have two coins, coin A and coin B, and flip them each 10 times. Let E be the event that every time coin A comes up heads, so does coin B. Find P(E).

hint: use conditional probability
Answer #1

probability that a coin comes up head, p = 0.5

tossing a coin 10 times, each occurance is independent of the other, hence probability of getting all heads on any coin is 0.5^10

Hence, P(E) = 0.5^10 * 0.5^10 = 0.5^20